[figure: 9.7 cm]

SPIN AND CHARGE FLUCTUATIONS IN THE DYNAMICAL

MEAN FIELD THEORY OF STRONGLY CORRELATED ELECTRON SYSTEMS

A.M. Shvaika Institute for Condensed Matter Physics Natl. Acad. Sci. Ukr. 1 Svientsitskii Str., UA-79011 Lviv, Ukraine

The contribution of the local spin and charge fluctuations in the self-energy and densityof-states are considered for strongly correlated electron systems within the strong coupling analytical approach [1,2] in the Dynamical Mean Field Theory (DMFT). The Luttinger-Ward functional Φ, which includes as Hartree-Fock contributions as the selfconsistent renormalization (SCR) of bosonic excitations, is constructed that allows to perform all calculations in the thermodynamically consistent way. It is shown for the Hubbard model that spin fluctuations give the main contribution close to half filling in the non-Fermi liquid regime while for small electron (n ≪ 1) and hole (2 − n ≪ 1) concentrations in the Fermi liquid regime the contributions from the charge fluctuations are essential.

[1] A.M. Shvaika, Phys. Rev. B 62, 2358 (2000)

[2] A.M. Shvaika, Acta Physica Polonica B 32, 3415 (2001)

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : A.M. Shvaika

Address for correspondence : Institute for Condensed Matter Physics Natl. Acad. Sci. Ukr. 1 Svientsitskii Str., UA-79011 Lviv, Ukraine

Email address : ashv@icmp.lviv.ua